# AsymFormer: Asymmetrical Cross-Modal Representation Learning for Mobile Platform Real-Time RGB-D Semantic Segmentation

### v5

@@ Abstract
In the realm of robotic intelligence, achieving efficient and precise RGB-D semantic segmentation is a key cornerstone. State-of-the-art multimodal semantic segmentation methods, primarily rooted in symmetrical skeleton networks, find it challenging to harmonize computational efficiency and precision. In this work, we propose AsymFormer, a novel network for real-time RGB-D semantic segmentation, which targets the minimization of superfluous parameters by optimizing the distribution of computational resources and introduces an asymmetrical backbone to allow for the effective fusion of multimodal features. Furthermore, we explore techniques to bolster network accuracy by redefining feature selection and extracting multi-modal self-similarity features without a substantial increase in the parameter count, thereby ensuring real-time execution on robotic platforms. Additionally, a Local Attention-Guided Feature Selection (LAFS) module is used to selectively fuse features from different modalities by leveraging their dependencies. Subsequently, a Cross-Modal Attention-Guided Feature Correlation Embedding (CMA) module is introduced to further extract cross-modal representations. This method is evaluated on NYUv2 and SUNRGBD datasets, with AsymFormer demonstrating competitive results with 54.1% mIoU on NYUv2 and 49.1% mIoU on SUNRGBD. Notably, AsymFormer achieves an inference speed of 65 FPS and after implementing mixed precision quantization, it attains an impressive inference speed of 79 FPS on RTX3090. This significantly outperforms existing multi-modal methods, thereby demonstrating that AsymFormer can strike a balance between high accuracy and efficiency for RGB-D semantic segmentation. Code: https://github.com/Fourier7754/AsymFormer
I INTRODUCTION
Real-time semantic segmentation methods are important for robots as they rely on obtaining semantic information in real-time to support navigation and task decision-making. Current Real-Time semantic segmentation networks achieve competitive results and reach over 100 FPS inference speed on simple outdoor environments (Cityscapes 19 classes) [3, 4, 2, 1]. However, these networks perform poorly on complex indoor environments, such as NYU Depth v2 (40 classes) [5] and SUNRGBD (37 classes) [6]. Several studies have explored how to improve indoor scene semantic segmentation performance by integrating RGB-D information [7, 8, 9, 10], which is widely available from RGB-D sensors on indoor robots. RGB-D information consists of RGB (color, texture and shape) and Depth (boundaries and relative location) features, which are somewhat complementary [7, 8].
At present, RGB-D semantic segmentation methods have achieved state-of-the-art performance by employing dual-branch backbone for feature extraction and attention mechanisms for feature selection [8]. However, there is a dearth of discourse on whether these operations can be executed with less redundancy and higher efficiency. Primarily, most methodologies utilize two symmetrical, computationally intensive backbones, which effectively doubles the overall computational complexity [12]. Secondly, numerous methods concentrate on designing complex feature fusion modules to enhance accuracy, but do not exhibit significant advantages than some methods that employ simpler attention modules like ESANet [9], which indicates that as more computational resources are invested, the incremental improvement in accuracy decreases. Lastly, the use of cross-attention modules to jointly utilize multi-modal features has been demonstrated to be effective in improving network accuracy [8, 28]. However, most methods do not discuss how to implement cross-attention more efficiently. Therefore, the reduction of redundancy in dual-branch networks and the construction of more efficient attention modules remain significant topics for discussion.
To address this issue, this study suggests a balanced approach between computational efficiency and feature richness in multi-modal representation learning. RGB images are found to contain more information than Depth images, thus it is not necessary to use the same size backbone for extracting Depth information. The performance improvements achieved by using attention mechanisms for various tasks come with minimal increase in computational cost. While existing modules like SE [34] are efficient for feature selection, they lack the ability to model correlations in spatial dimensions. Increasing parameters can facilitate spatial dimension feature selection, but careful design is needed to maintain computational efficiency. For feature extraction, the commonly used Multi-Head Self Attention (MHSA) [22] is effective for single-modal features despite its small parameter count. Extending this efficient feature extraction to multiple modalities is feasible, but consideration should be given to feature fusion while maintaining computational efficiency.
@@ I INTRODUCTION
This paper introduces AsymFormer, a high-performance real-time network for RGB-D integration semantic segmentation that employs an asymmetrical backbone design. This includes a larger parameter backbone for RGB features and a smaller backbone for the Depth branch. Regarding framework selection, Vision Transformer are always performance better, but slower than same computational complexity CNNs due to a lack of hardware optimization [40]. In order to speed up the main branch, this paper employ a hardware-friendly CNN [16] for the RGB branch and a Transformer [17] with fewer parameters but similar performance for the Depth branch to further compress the parameters. This paper also discuss feature selection modeling and construct the Local Attention-Guided Feature Selection (LASF) module. This module computes the spatial-channel attention weights in parallel to enhance inference speed. Moreover, this paper establishes a novel approach to model spatial attention. This module can estimate the differences in features from different modalities by utilizing learnable channel weights and calculating spatial attention weights for each pixel. Additionally, a Cross-Modal Attention (CMA) module is introduced to embed cross-modal integration learning information into pixel-wise fused features. Finally, we employ a lightweight MLP-Decoder[17] to decode semantic information from shallow features.
@@
This paper evaluates AsymFormer on two classic indoor scene semantic segmentation datasets: NYUv2 and SUNRGBD. Meanwhile, the inference speed test is also preformed on Nvidia RTX 3090 platform. The AsymFormer achieves 54.1% mIoU on NYUv2 and 49.1 mIoU on SUNRGBD, with 65 FPS inference speed. After performing mixed precision quantization on the model using the TensorRT framework, AsymFormer achieves an impressive inference speed of 79 FPS on RTX 3090, which significantly surpasses existing methods. Our experiments highlight AsymFormer’s ability to acquire high accuracy and efficiency at the same time. Our main contributions are summarized as follows:
@@
II Related Works
RGB-D Representation Learning. One of the earliest works on RGB-D semantic segmentation, FCN [35], treated RGB-D information as a single input and processed it with a single backbone. However, subsequent works have recognized the need to extract features from RGB and Depth information separately, as they have different properties. Therefore, most of them have adopted two symmetric backbones for RGB and Depth feature extraction [36, 37, 9, 11, 27, 8]. Further, the ACNet [25] proposed a three-stream backbone to process RGB, Depth and Fusion features independently. Moreover, some works have considered the variability of RGB-D features and used asymmetric backbones for their feature extraction. For example, TSNet [39] used ResNet [31] for RGB feature extraction and VGG (without residual connections) [38] for Depth feature extraction. The PSCNet [12] also used an asymmetric backbone and reduced the redundant computational cost by cutting the parameters of the Depth branch.
RGB-D feature fusion. The design of RGB-D backbones for feature extraction has been mostly based on two or three stream architectures [8]. However, the performance of different frameworks depends largely on how they fuse RGB and Depth features. Some early works, such as FuseNet[36] and RedNet [37], fused RGB and Depth feature maps pixel-wise in the backbone. Later works, such as ACNet [25], ESANet [9] and EMSANet[11], proposed channel attention to select features from different channels, as RGB and Depth feature maps may not align well on the corresponding channels. PSCNet [12] further extended channel attention to both spatial and channel directions and achieved better performance. Recently, more complex models have been proposed to exploit cross-modal information and select features for RGB-D fusion. For example, SAGate [27] proposed a gated attention mechanism that can leverage cross-modal information for feature selection. CANet [28] extended non-local attention [20] to cross-modal semantic information and achieved significant improvement. CMX [8] extended SA-Gate to spatial and channel directions and proposed a novel cross-modal attention with global receptive field, which can enhance cross-modal complementarity using the similarity of two modalities. However, integrating cross-modal information and learning cross-modal similarity is still an open question in vision tasks.
III Method
@@
III-A Framework Overview
This paper proposes AsymFormer, a novel network for RGB-D integration semantic segmentation. The network employs a ConvNext [16] based backbone to extract features from RGB images and a Mix-Transformer [17] based backbone to process RGB-D fused features. We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights. We integrated modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters. The overall framework of AsymFormer is illustrated in Figure 1.
III-B Local Attention-Guided Feature Selection
Attention mechanisms have been shown to be effective in selecting complementary features from RGB and Depth features, thus improving efficiency and performance[25, 1, 12, 8]. Current spatial attention methods extract global information using pixel maximum or average values, neglecting the differences in channels from multiply modalities. In this work, we extend the previous spatial attention by proposing a more flexible spatial attention mechanism. The core design of LAFS is a trainable spatial attention weight. This weight is learned through an MLP in a similar manner to SE [34].
In specific, Figure 1 (b)-1 illustrates the details of LAFS. The LAFS receives input features of the concatenation of RGB feature $RGB$ with dimension $(C_{1},H,W)$ and current layer Depth feature $Depth$ with dimension $(C_{2},H,W)$.
For the computation of the attention weight, we first extract global information vectors $Avg$ of each feature map by adaptive AvgPool. For the channel attention, we process the global information by a feed-forward network with a squeeze-and-excitation structure.
$W_{C}=\text{Sigmoid}(\text{FFN}(Avg))$ (1)
For the spatial attention, we process the global information $Avg$ by another feed-forward network with a squeeze-and-excitation structure. The output vector is split into $R_{Max}$ and $R_{Avg}$, which represent two different descriptions of pixel spatial-similarity.
$R_{Avg}=\text{FFN}(Avg)$ (2)
Furthermore, the global spatial information $I_{S}$ of feature map can be extracted by calculating the inner product similarity from $R_{Avg}$ and the $Input$ (In computation, this is equivalent to the weighted sum of the features from each channel). Then, the spatial attention weight $W_{S}$ is computed by sigmoid normalization.
@@ III-B Local Attention-Guided Feature Selection
[…]
$Output=W_{C}\times W_{S}\times Input$ (4)
III-C Cross Modal Attention-Guided Feature Correlation Embedding
In this paper, we introduce a novel cross-modal attention (CMA) module that incorporates cross-modal self-similarity information into pixel-wise fused features. Current MHSA [21] is limited to learning the self-similarity of a single modality, whereas in multi-modal processing, our goal is to jointly use information from multiple modalities for representation learning. To achieve this, CMA is proposed to learn the self-similarity of multiple modalities. The key insight of CMA is independently learning the self-similarity of two modalities and embeds the sum of the results into the fusion feature, which is different than standard MHSA in theoretical. Specifically, for each pixel $(i,j)$ in feature map, we separately embed its RGB and Depth feature:
$Key_{RGB}=(Kr_{1,i,j},Kr_{2,i,j},...,Kr_{N,i,j})$ (5)
$Query_{RGB}=(Qr_{1,i,j},Qr_{2,i,j},...,Qr_{N,i,j})$
$Key_{Depth}=(Kd_{1,i,j},Kd_{2,i,j},...,Kd_{N,i,j})$
$Query_{Depth}=(Qd_{1,i,j},Qd_{2,i,j},...,Kd_{N,i,j})$
The cross-modal similarity of pixel $(i_{0},j_{0})$ and other pixel in feature map are defined as:
$W(i,j)=\sum_{n=1}^{N}(Kr_{n,i,j}\cdot Qr_{n,i_{0},j_{0}})+\sum_{n=1}^{N}(Kd_{n,i,j}\cdot Qd_{n,i_{0},j_{0}})\$ (6)
Moreover, the CMA embeds cross-modal similarity into the fused feature $Value$ with same operation as[21, 20, 22]. Following above idea, the CMA first embedding $Key$, $Query$ and $Value$ independently for three modalities features. Then, we extend self-attention to the multimodal case, which contain two design: 1. Information mixing and redistribution of $Key$ and $Query$; 2. Learning two different different representation subspaces for the same position.
Independent linear embedding: According to our definition, the $Key$, $Query$ and $Value$ should be embedded independently. Assuming the input $RGB$ feature has dimension ${}^{(}C_{0},H,W)$; the Enhanced Depth feature $Depth_{En.}$ has dimension ${}^{(}C_{1},H,W)$; and the previous fused feature $Fus_{1}$ has dimension $(C_{2},H,W)$. We first normalize the RGB feature and Depth feature. Then, we use six independent convolution layers to generate self-attention keys and queries from RGB and Depth feature:
$RGB_{emb}=\left\{\begin{array}[]{c}K_{RGB}=\text{Conv}_{f\times f}(RGB)\\ Q_{RGB}=\text{Conv}_{1\times 1}(RGB)\\ \end{array}\right.$ (7)
$Depth_{emb}=\left\{\begin{array}[]{c}K_{Depth}=\text{Conv}_{f\times f}(Depth_{En.})\\ Q_{Depth}=\text{Conv}_{1\times 1}(Depth_{En.})\\ \end{array}\right.$ (8)
$Fused_{emb}=\left\{\begin{array}[]{c}V=\text{Conv}_{f\times f}(Fused_{1})\\ V_{1},V2=\text{Split}(V,2)\\ \end{array}\right.$ (9)
The dimensions of $K_{RGB}$ and $K_{Depth}$ are $(\frac{C_{1}}{4},\frac{H}{f},\frac{W}{f})$, where $f$ is the down-sample rate and equals the convolution kernel-size and stride. This reduces the feature map resolution to $\frac{1}{f}$. The dimensions of $Q_{RGB}$ and $Q_{Depth}$ are $(\frac{C_{1}}{4},H,W)$. The $V_{1}$ and $V_{2}$ is reshaped into dimension $(\frac{C_{2}}{4},\frac{H\times W}{f^{2}})$.
Information mixing and redistribution: We concatenate the cross-modal keys and queries to obtain the $Key$ and $Query$:
@@ III-C Cross Modal Attention-Guided Feature Correlation Embedding
The first $\frac{C_{2}}{4}$ channels of $Key$ and $Query$ contain RGB information. Similarly, the second half of the $Key$ and $Query$ channels contain Depth information. To obtain the $K1$, $K2$ and $Q1$, $Q2$ for integrating RGB-D information computing the two feature subspaces $W_{1}$, $W_{2}$, we perform a channel shuffle of $Key$ and $Query$ and split them into required vector:
$K_{1},K_{2}=\text{Split}(\text{Shuffle}(Key),2)$ (11)
$Q_{1},Q_{2}=\text{Split}(\text{Shuffle}(Query),2)$
Representation subspaces learning: The $Q_{1}$, $Q_{2}$ are reshaped into dimension $(\frac{C_{1}}{4},H\times W)$ and the $K_{1}$, $k_{2}$ are reshaped into dimension $(\frac{C_{1}}{4},\frac{H\times W}{f^{2}})$. Then, the two representation subspaces $W_{1}$ and $W_{2}$ are computed by dot product of $K$ and $Q$ respectively:

### v7

@@ Abstract
Understanding indoor scenes is crucial for urban studies. Considering the dynamic nature of indoor environments, effective semantic segmentation requires both real-time operation and high accuracy.To address this, we propose AsymFormer, a novel network that improves real-time semantic segmentation accuracy using RGB-D multi-modal information without substantially increasing network complexity. AsymFormer uses an asymmetrical backbone for multimodal feature extraction, reducing redundant parameters by optimizing computational resource distribution. To fuse asymmetric multimodal features, a Local Attention-Guided Feature Selection (LAFS) module is used to selectively fuse features from different modalities by leveraging their dependencies. Subsequently, a Cross-Modal Attention-Guided Feature Correlation Embedding (CMA) module is introduced to further extract cross-modal representations. The AsymFormer demonstrates competitive results with 54.1% mIoU on NYUv2 and 49.1% mIoU on SUNRGBD. Notably, AsymFormer achieves an inference speed of 65 FPS (79 FPS after implementing mixed precision quantization) on RTX3090, demonstrating that AsymFormer can strike a balance between high accuracy and efficiency. Code: https://github.com/Fourier7754/AsymFormer
111$\nmid$ indicates equal contribution.222* corresponding author.
Figure 1: The AsymFormer has 33.0 million parameters and 36.0 GFLOPs computational cost, and it can achieve 65 FPS inference speed on RTX 3090, 54.1% mIoU on NYUv2.
1 Introduction
Indoor scenes are essential to urban environments. Current urban studies necessitate understanding indoor scene semantic information for tasks like emergency evacuation [38], robotic navigation [36], and virtual reality [43]. The dynamic nature of indoor scenes demands perception of environmental changes in real time for tasks such as emergency evacuation [38], stressing the importance of algorithms’ real-time capabilities. Existing real-time semantic segmentation methods often falter when applied to the complex semantic information of indoor scenes, usually requiring a sacrifice in accuracy [40]. This often requires a trade-off between inference speed and increased network complexity to achieve adequate segmentation accuracy indoors. Consequently, a critical research question is how to improve the accuracy of semantic segmentation in indoor environments without substantially increasing complexity, while ensuring real-time performance.
Aside from increasing the complexity of the network, introducing additional information, such as RGB-D data, is also an effective way to improve the accuracy of semantic segmentation networks. RGB-D cameras are widely used devices for indoor information acquisition. RGB-D information consists of RGB (color, texture and shape) and Depth (boundaries and relative location) features, which are somewhat complementary [42]. Several studies have explored how to improve indoor scene semantic segmentation performance by integrating RGB-D information [42, 26, 50].
Existing research has explored the implementation of attention mechanisms to extract valuable information from RGB-D features without significantly increasing computational complexity. However, due to the additional feature extraction branch for depth features and the lack of discussion on how to allocate computational resources based on feature importance, these methods often introduce a substantial amount of redundant parameters, significantly reducing their computational efficiency [9].
@@ 1 Introduction
To address this issue, this paper introduces AsymFormer, a high-performance real-time network for RGB-D semantic segmentation that employs an asymmetric backbone design. This includes a larger parameter backbone for important RGB features and a smaller backbone for the Depth branch. Regarding framework selection, at the same computational complexity, Transformer often achieves higher accuracy but has a slower inference speed compared to CNN [19]. In order to speed up the main branch, this paper employ a hardware-friendly CNN [21] for the RGB branch and a light-weight Transformer [35] for the Depth branch to further compress the parameters. Considering the differences between different modal representation, to effectively select and fuse asymmetric features, this paper proposes a learnable method for feature information compression and constructs a Local Attention Guided Feature Selection (LAFS) module. Additionally, a Cross-Modal Attention (CMA) module is introduced to embed cross-modal information into pixel-wise fused features. Finally, we employ a lightweight MLP-Decoder[35] to decode semantic information from shallow features.
@@
This paper evaluates AsymFormer on two classic indoor scene semantic segmentation datasets: NYUv2 and SUNRGBD. Meanwhile, the inference speed test is also preformed on Nvidia RTX 3090 platform. The AsymFormer achieves 54.1% mIoU on NYUv2 and 49.1 mIoU on SUNRGBD, with 65 FPS inference speed (79 FPS with mixed precision quantization using the TensorRT). Our experiments highlight AsymFormer’s ability to acquire high accuracy and efficiency at the same time. The main contributions are summarized as follows:
@@
2 Related Works
@@
2.1 Indoor Scene Understanding
Current research in indoor scene understanding leverages diverse data sources, including RGB images [2], RGB-D images [42], point clouds [51], and mesh data [45]. Each of these sources provides unique insights and benefits for analyzing and interpreting indoor environments. For instance, RGB images are accessible and straightforward for visual representations, whereas point clouds and meshes offer intricate 3D spatial data. However, when considering the simultaneous use of both 2D and 3D information to optimize efficiency and effectiveness in scene understanding, RGB-D images emerge as the optimal choice.
2.2 RGB-D Representation Learning
One of the earliest works on RGB-D semantic segmentation, FCN [22], treated RGB-D information as a single input and processed it with a single backbone. However, subsequent works have recognized the need to extract features from RGB and Depth information separately, as they have different properties. Therefore, most of them have adopted two symmetric backbones for RGB and Depth feature extraction [26, 27, 5, 42]. Primarily, asymmetric backbones doubles the overall computational complexity [9]. However, it is generally observed that for semantic segmentation, RGB information typically plays a more prominent role than Depth information, as indicated by [9]. Using a asymmetric backbone for feature extraction will obviously lead to redundant parameters on the less important side, reducing the efficiency of the network.
2.3 RGB-D feature fusion
The performance and efficiency of different frameworks depends largely on how they fuse RGB and Depth features. Some early works, such as RedNet [17], fused RGB and Depth feature maps pixel-wise in the backbone. Later, ESANet series [26, 27] proposed channel attention to select features from different channels, as RGB and Depth feature maps may not align well on the corresponding channels. PSCNet [9] further extended channel attention to both spatial and channel directions and achieved better performance. Recently, more complex models have been proposed to exploit cross-modal information and select features for RGB-D fusion. For example, SAGate [5] proposed a gated attention mechanism that can leverage cross-modal information for feature selection. CANet [47] extended non-local attention [32] to cross-modal semantic information and achieved significant improvement. CMX [42] extended SA-Gate to spatial and channel directions and proposed a novel cross-modal attention with global receptive field. However, integrating cross-modal information and learning cross-modal similarity is still an open question in vision tasks.
3 Method
3.1 Framework Overview
This paper develops a high-accuracy real-time semantic segmentation method, AsymFormer, which leverages an asymmetric backbone network to mine multimodal representations. A hardware friendly convolution network ConvNext [21] is used for RGB feature extraction and a light-weight Mix-Transformer [35]is used for processing Depth feature. To effectively fuse RGB-D features, the study introduces a Local Attention Guided Multimodal Feature Selection (LAFS) module, which uses learnable strategy to extract global information and selects multimodal features in both spatial and channel dimensions. Moreover, to further mine the information contained in multimodal features, the study embeds the information contained in multimodal features through a novel Cross-Modal Attention (CMA) module. The overall framework of AsymFormer is shown in Fig.2.
3.2 Local Attention Guided Feature Selection
Studies have shown that attention mechanisms can effectively select complementary features from RGB and Depth features, thereby improving the efficiency of feature extraction and the performance of semantic segmentation [16, 40, 9, 42]. However, existing attention mechanisms usually adopt a non-learnable, fixed strategy for feature information compression [34], which may overlook the differences among various modal features, causing insufficient use of information.
Figure 3: LAFS.
In response to this issue, this paper proposes a learnable method for feature information compression and constructs a Local Attention Guided Feature Selection (LAFS) module. The core design of LAFS is a learnable method for spatial information compression. Inspired by SE [14] attention, LAFS abandons traditional fixed strategies and employs a feedforward neural network to learn a set of dynamic spatial information compression rules.
The Fig. 3 demonstrates details of the LAFS module. The input features to the LAFS module are the tensor concatenation of $RGB$ and $Depth$ features. For computing attention weights, the study first extracts a global information vector $Avg$ through Adaptive Average Pooling. For channel attention, its structure is the same as SE [14].
On spatial attention computation, which is the primary improvement, the LAFS first processes global information $Avg$ through another feedforward neural network with a squeeze-excitation structure, producing the output vector $R_{Avg}$, representing pixel space similarity descriptions. Then, by computing the dot product similarity between $R_{Avg}$ and input feature map, the global spatial information $I_{S}$ is extracted (this equates to a weighted sum of the channel features, dynamically varying with $Avg$ changes). Subsequently, spatial attention weights $W_{S}$ are calculated through sigmoid normalization.
@@ 3.2 Local Attention Guided Feature Selection
Here, all results are divided by $C^{2}$ to prevent sigmoid overflow. Finally, input features are selected based on the pixel-wise multiplication of feature maps with spatial and channel attention weights.
3.3 Cross-Attention Guided Feature Embedding
Aside from selecting existing features, it is necessary to mine new information from fused features. Existing Multi-Head Self-Attention (MHSA) in Transformer [31] is limited to learning self-similarity within a single modality, whereas mining information from multi-modal jointly is a new goal in representation learning. In this paper, a new Cross-Modal Attention (CMA) module is constructed. The key to CMA is defining cross-modal self-similarity using a linear sum, and embedding its result into the fused features.
3.3.1 Definition of Cross-Modal Self-Similarity:
Suppose $RGB$ and $Depth$ features are seamlessly embedded into $Key$ and $Query$, then for a pixel $(i_{0},j_{0})$, its cross-modal self-similarity with other pixels $(i,j)$ can be defined as:
$W(i,j)=\sum_{n=1}^{N}(Kr_{n,i,j}\cdot Qr_{n,i_{0},j_{0}})+\sum_{n=1}^{N}(Kd_{n,i,j}\cdot Qd_{n,i_{0},j_{0}})$ (2)
where $Kr_{n,i,j}$ represents the nth feature value of the pixel $Key_{RGB}$, and $Qr_{n,i,j}$ represents the nth feature value of the pixel $Query_{RGB}$. Similarly, $Kd_{1,i,j}$ and $Qd_{1,i,j}$ represent the nth feature value of the pixel $Key_{Depth}$ and $Query_{Depth}$ respectively.
3.3.2 Feature Embedding
The CMA has three input features, namely $RGB$ features, $Depth$ features, and the fused features $Fused$ selected by LAFS. In the calculation process of CMA, the first step is to embed the input features $RGB$, $Depth$, and $Fused$ into vector space.
Figure 4: Feature Embedding.
Simultaneously, the output $Value$ split into two independent vectors $V_{1}$ and $V_{2}$ in the channel direction.
3.3.3 Splitting and Mixing of Multimodal Information
To learn features from multiple subspaces, the embedded features $K_{RGB}$, $K_{Depth}$, $Q_{RGB}$, and $Q_{Depth}$ are split into two independent vectors.The study first concatenates $Key_{RGB+Depth}$ with $Query_{RGB+Depth}$ in the channel direction, obtaining $Key$ and $Query$:
@@ 3.3.3 Splitting and Mixing of Multimodal Information
It can be noted that if the vectors are directly split and self-similarity is calculated in different subspaces, it might result in the inability to simultaneously include RGB and Depth features in different subspaces. Therefore, we introduce a Shuffle mechanism to ensure each $K1$, $K2$, $Q1$, and $Q2$ contain information from both modalities. As shown in Figure 4.
Figure 5: Splitting and Mixing of Multimodal Information.
After channel shuffle, CMA can ensure that both RGB and Depth information can be utilized simultaneously when calculating self-similarity.
3.3.4 Representation Learning in Multiple Subspaces
Finally, CMA calculates the cross-modal similarity and embeds the calculation result into the fused feature $Value$, similar to operations in [31, 32, 8]. Specifically, the study computes the dot products $K_{1}$, $Q_{1}$ and $K_{2}$, $Q_{2}$ to derive two representations of subspaces $W_{1}$ and $W_{2}$: